OpenStudy (anonymous):

is xy-3x=y+3 a function

OpenStudy (tkhunny):

xy - y = 3 + 3x y(x-1) = 3 + 3x \(y = \dfrac{3\cdot(1+x)}{x-1}\) Did we learn anything?

OpenStudy (anonymous):

Yes thank you so much

OpenStudy (tkhunny):

What was your answer?